# Cobordism Obstructions to Complex Sections II: Torsion Obstructions

### v1

@@ 3. Computing the torsion obstruction
Fix prime $p>2$. Consider the cobar complexes for $\overline{\mathbf{MTU}}(d)$. We will consider the cobar complex as the $E_{1}$ page of the Adams-Novikov spectral sequence. Suppose we have an element $\beta\in E_{2}^{s,t}(\overline{\mathbf{MTU}}(d))$. This element is represented by $\sum_{k}\mu_{k}\otimes x_{k}\in\Gamma^{\otimes s}\otimes BP_{*}(\mathbf{MTU}(d))$ such that $\sum_{k}\mu_{k}\otimes x_{k}\in\textnormal{Ker}(d_{1})$. Suppose $t$ is such that $E^{s,r}_{2}(\Sigma^{\infty}S^{0})=0$ for all $r\leq t-(2d+2)$. In particular for $s=2$, this is equivalent to supposing $t<2(p^{2}-p+d+1)$. The element $\sum_{k}\mu_{k}\otimes x_{k}$ has a term with maximal degree $x_{k}$. We will call this term $\mu_{i}\otimes x_{i}$. Since $\sum_{k}\mu_{k}\otimes x_{k}\in\textnormal{Ker}(d_{1})$, we see that $d_{1}(\mu_{i})=0$. We know $\mu_{i}$ has degree less than or equal to $t-(2d+2)$ because $x_{i}$ has degree greater than or equal to $(2d+2)$. Thus $\mu_{i}\otimes 1$ is in the image of $\sum_{j}\tilde{\mu}_{j}\otimes\nu_{j}$ where $\tilde{\mu}_{j}\in\overline{\Gamma}^{\otimes s-1}$ and $\nu_{j}\in\pi_{*}(BP)$ because the corresponding element must vanish in $E_{2}(S^{0})$. Then, we observe that $d_{1}(\sum_{j}\tilde{\mu}_{j}\otimes\nu_{j}x_{i})=\mu_{i}\otimes x_{i}$. Thus, $\sum_{k\neq i}\mu_{k}\otimes x_{k}$ also represents $\beta$. Repeating this process, $\beta$ can be represented by 00, and thus $\beta=0$. (We choose the letter $\beta$ suggestively, as the first potentially non zero $\beta$ corresponds to $\beta_{1}$ constructed in Definition 1.3.14 [11]). By Proposition 3.8, $E_{2}^{3,t}(\mathbf{MTU}(d))=0$ and $t<2(p^{2}-p+d+1)$. Thus, $d_{3}:E_{3}^{0,t}\rightarrow E_{3}^{3,t+2}$ is trivial for $t<2(p^{2}-p+d)$. Similarly, the higher differentials must also vanish. By Proposition 3.6, the torsion obstruction to $r<p^{2}-p$ complex sections vanishes. ∎
@@
We now focus on the case $p=2$. Thus for the rest of the paper all homotopy groups are assumed to be 2-primary. We will need to compute the homotopy groups of $\overline{\mathbf{MTU}}(d)$. To do so we remind ourselves of the Adams-Novikov spectral sequence for the sphere spectrum: [14]
@@
Once again, if $d$ is even, then the cokernel of $d_{1}:E_{1}^{1,2d+6}\rightarrow E_{1}^{2,2d+6}$ is $\mathbb{Z}/2$, and if $d$ is odd, the cokernel is zero. We omit the details, but for similar reasons, $d_{1}:E_{1}^{2,2d+8}\rightarrow E_{1}^{3,2d+8}$ has cokernels $\mathbb{Z}/2$ or 00 when $d$ is even or odd respectively.
We now compute the image of $d_{1}:E_{1}^{0,2d+6}\rightarrow E_{1}^{1,2d+6}$. We will write this as a table for brevity.
$\begin{array}[]{c|c|c|c|c}&t_{1}\otimes B_{2}B_{1}^{d},&t_{1}\otimes B_{1}^{d+2}&t_{1}\otimes v_{1}B_{1}^{d+1}&t_{1}^{2}\otimes B_{1}^{d+1}\\ B_{3}B_{1}^{d}&3&0&-2&5\\ B_{2}^{2}B_{1}^{d-1}&4&0&0&4\\ B_{2}B_{1}^{d+1}&(d+1)&2&0&2d+3\\ B_{1}^{d+3}&0&(d+3)&0&\frac{1}{2}(d+3)(d+2)\\ v_{1}B_{2}B_{1}^{d}&-2&0&2&-4\\ v_{1}B_{1}^{d+2}&0&-2&(d+2)&2(d+2)\\ v_{1}^{2}B_{1}^{d+1}&0&0&-4&4\\ \end{array}$
The kernel of $d_{1}:E_{1}^{0,2d+6}\rightarrow E_{1}^{1,2d+6}$ is generated by elements where $02a_{1}+(d+2)a_{2}-2(a_{3}+a_{4})=$ 0:
@@
Thus, we have two cases for $E_{2}^{1,2d+4}$: if $d$ is even, $E_{2}^{1,2d+4}=\mathbb{Z}/2$ and if $d$ is odd, $E_{2}^{1,2d+4}=0$. Similarly we can compute $d_{1}:E_{1}^{1,2d+8}\rightarrow E_{1}^{2,2d+8}$. We find that $E_{2}^{2,2d+6}$ is always zero.
$\begin{array}[]{c|c|c|c|c|c}&t_{1}\otimes t_{1}\otimes B_{2}B_{1}^{d},&t_{1}\otimes t_{1}\otimes B_{1}^{d+2}&t_{1}\otimes t_{1}\otimes v_{1}B_{1}^{d+1}&t_{1}\otimes t_{1}^{2}\otimes B_{1}^{d+1}&t_{1}^{2}\otimes t_{1}\otimes B_{1}^{d+1}\\ t_{1}\otimes B_{3}B_{1}^{d}&3&0&-2&5&0\\ t_{1}\otimes B_{2}^{2}B_{1}^{d-1}&4&0&0&4&0\\ t_{1}\otimes B_{2}B_{1}^{d+1}&(d+1)&2&0&2d+3&0\\ t_{1}\otimes B_{1}^{d+3}&0&(d+3)&0&\frac{1}{2}(d+3)(d+2)&0\\ t_{1}\otimes v_{1}B_{2}B_{1}^{d}&-2&0&2&-4&0\\ t_{1}\otimes v_{1}B_{1}^{d+2}&0&-2&(d+2)&2(d+2)&0\\ t_{1}\otimes v_{1}^{2}B_{1}^{d+1}&0&0&-4&4&0\\ t_{1}^{2}\otimes B_{2}B_{1}^{d}&-2&0&0&0&2\\ t_{1}^{2}\otimes B_{1}^{d+2}&0&-2&0&0&(d+2)\\ t_{1}^{2}\otimes v_{1}B_{1}^{d+1}&0&0&-2&0&-2\\ t_{2}\otimes B_{1}^{d+1}&0&0&-1&3&2\\ t_{1}^{3}\otimes B_{1}^{d+1}&0&0&0&3&3\\ \end{array}$
Then, in the relevant range the Adams chart for $d$ even is:
@@
The non-trivial $d_{3}$ differential can be observed by looking at the map $S^{\infty+2d}\rightarrow\overline{\mathbf{MTU}}(d)$ given by the generator of $\pi_{2d}(\overline{\mathbf{MTU}}(d))$. There may be an additional differential killing $E_{2}^{2,2d+8}$ but since this element is in odd degree, we will not need to know it. The Adams-Novikov spectral sequence for $d$ odd is trivial in this range. ∎
@@
Since $\pi_{2d+6}(\overline{\mathbf{MTU}}(d))$ is free abelian, there is no possible target for a torsion obstruction to $3$ complex sections. Similarly, there is no possible target for a torsion obstruction to $2$ complex sections when $d$ is odd. In the case $d$ is even, there could be an obstruction to $2$ complex sections. However, we recall from [8] that $S^{\infty+2d}$ splits off $MTU(d)$ for $d$ even. The element in $E_{2}^{3,2d+3}$ comes from this split $S^{\infty+2d}$ (specifically, it can be identified with $h_{1}^{3}$), and so cannot be the target of a differential. ∎
@@
We know that $\pi_{6}^{s}(BU(d-3))$ is torsion free. The torsion element in $\pi_{6}^{s}(S^{0})\cong h_{2}\cdot h_{2}$. The map $\pi_{2d}^{s}(S^{2d-6})\rightarrow\pi_{2d}(MTU(d-3))$ is a $\pi_{*}^{s}(S^{0})$ module map. In particular, it must take $h_{2}$ to 00 if $d$ is odd and we conclude that $h_{2}\cdot h_{2}\mapsto h_{2}\cdot 0=0$. Thus the torsion obstruction to $M$ having $4$ complex sections vanishes. ∎
@@
There are no non-zero differentials originating in the bottom row of the Adams-Novikov spectral sequence for $MTU(1)\cong\Sigma^{\infty-2}\mathbb{C}P^{\infty}$.
@@
Then for any set of multiplicative generators $M^{2d}\in\Omega_{2d}^{U}$, we can write $N^{2d}=\tilde{a}_{d}M^{2d}+\textnormal{decomposables}$

### v2

@@ 3. Computing the torsion obstruction
Fix prime $p>2$. Consider the cobar complex for $\overline{\mathbf{MTU}}(d)$. Suppose we have some element $\beta\in E_{2}^{2,t}(\overline{\mathbf{MTU}}(d))$ for $t<2(p^{2}-p+d+1)$ This element is represented by the tensor sum $\sum_{k}\mu_{k}\otimes x_{k}\in\Gamma^{\otimes s}\otimes BP_{*}(\mathbf{MTU}(d))$ such that $\sum_{k}\mu_{k}\otimes x_{k}\in\textnormal{Ker}(d_{1})$. The element $\sum_{k}\mu_{k}\otimes x_{k}$ has a term with maximal degree $x_{k}$. We will call this term $\mu_{i}\otimes x_{i}$. Since the sum $\sum_{k}\mu_{k}\otimes x_{k}\in\textnormal{Ker}(d_{1})$, we see that $d_{1}(\mu_{i})$ must be 00. We know $\mu_{i}$ has degree less than $2(p^{2}-p)$ because $x_{i}$ has degree greater than or equal to $(2d+2)$. The element $\mu_{i}\otimes 1$ can be considered as an element in the cobar complex for the sphere spectrum, and since it is in $\textnormal{Ker}(d_{1})$, it represents an element in $E_{2}(S^{0})$. By [14][7.1], this element must be zero. Thus $\mu_{i}\otimes 1$ is in the image of $\sum_{j}\tilde{\mu}_{j}\otimes\nu_{j}$ where $\tilde{\mu}_{j}\in\overline{\Gamma}^{\otimes s-1}$ and $\nu_{j}\in\pi_{*}(BP)$. Then, we observe that $d_{1}(\sum_{j}\tilde{\mu}_{j}\otimes\nu_{j}x_{i})=\mu_{i}\otimes x_{i}$. Thus, $\sum_{k\neq i}\mu_{k}\otimes x_{k}$ also represents $\beta$. Repeating this process, $\beta$ can be represented by 00, and thus $\beta=0$. (We choose the letter $\beta$ suggestively, as the first potentially non zero $\beta$ corresponds to $\beta_{1}$ constructed in Definition 1.3.14 [11]). By Proposition 3.8, $E_{2}^{3,t}(\mathbf{MTU}(d))=0$ for $t<2(p^{2}-p+d+1)$. Thus, $d_{3}:E_{3}^{0,t}\rightarrow E_{3}^{3,t+2}$ is trivial for $t<2(p^{2}-p+d)$. The same argument shows that all higher differentials are zero. (In fact, the bound is even stronger.) By Proposition 3.6, the $p$-torsion obstruction to $r<p^{2}-p$ complex sections vanishes. ∎
Suppose $t$ is such that $E^{s,q}_{2}(\Sigma^{\infty}S^{0})=0$ for all $q\leq t$. In particular for $s=2$, this is equivalent to supposing $t<2(p^{2}-p)$.
@@
We now focus on the case $p=2$. For the rest of the paper all homotopy groups are assumed to be 2-primary. We will need to compute the homotopy groups of $\overline{\mathbf{MTU}}(d)$. To do so we remind ourselves of the Adams-Novikov spectral sequence for the sphere spectrum: [14]
@@
Once again, if $d$ is even, then the cokernel of $d_{1}:E_{1}^{1,2d+6}\rightarrow E_{1}^{2,2d+6}$ is $\mathbb{Z}/2$, and if $d$ is odd, the cokernel is zero. For similar reasons, $d_{1}:E_{1}^{2,2d+8}\rightarrow E_{1}^{3,2d+8}$ has cokernels $\mathbb{Z}/2$ or 00 when $d$ is even or odd respectively. We now compute the image of the map $d_{1}:E_{1}^{0,2d+6}\rightarrow E_{1}^{1,2d+6}$. We will write this as a table giving of coefficients in the basis $t_{1}\otimes B_{2}B_{1}^{d},t_{1}\otimes B_{1}^{d+2},t_{1}\otimes v_{1}B_{1}^{d+1},t_{1}^{2}\otimes B_{1}^{d+1}$
$\begin{array}[]{c|c|c|c|c}B_{3}B_{1}^{d}&3&0&-2&5\\ B_{2}^{2}B_{1}^{d-1}&4&0&0&4\\ B_{2}B_{1}^{d+1}&(d+1)&2&0&2d+3\\ B_{1}^{d+3}&0&(d+3)&0&\frac{1}{2}(d+3)(d+2)\\ v_{1}B_{2}B_{1}^{d}&-2&0&2&-4\\ v_{1}B_{1}^{d+2}&0&-2&(d+2)&2(d+2)\\ v_{1}^{2}B_{1}^{d+1}&0&0&-4&4\end{array}$
The kernel of $d_{1}:E_{1}^{0,2d+6}\rightarrow E_{1}^{1,2d+6}$ is generated by elements
@@
where $02a_{1}+(d+2)a_{2}-2(a_{3}+a_{4})=$ 0. Thus, we have two cases for $E_{2}^{1,2d+4}$: if $d$ is even, $E_{2}^{1,2d+4}=\mathbb{Z}/2$ and if $d$ is odd, $E_{2}^{1,2d+4}=0$. Similarly we can compute $d_{1}:E_{1}^{1,2d+8}\rightarrow E_{1}^{2,2d+8}$ in the basis: $t_{1}\otimes t_{1}\otimes B_{2}B_{1}^{d},t_{1}\otimes t_{1}\otimes B_{1}^{d+2},t_{1}\otimes t_{1}\otimes v_{1}B_{1}^{d+1},\\ t_{1}\otimes t_{1}^{2}\otimes B_{1}^{d+1},t_{1}^{2}\otimes t_{1}\otimes B_{1}^{d+1}$.
$\begin{array}[]{c|c|c|c|c|c}t_{1}\otimes B_{3}B_{1}^{d}&3&0&-2&5&0\\ t_{1}\otimes B_{2}^{2}B_{1}^{d-1}&4&0&0&4&0\\ t_{1}\otimes B_{2}B_{1}^{d+1}&(d+1)&2&0&2d+3&0\\ t_{1}\otimes B_{1}^{d+3}&0&(d+3)&0&\frac{1}{2}(d+3)(d+2)&0\\ t_{1}\otimes v_{1}B_{2}B_{1}^{d}&-2&0&2&-4&0\\ t_{1}\otimes v_{1}B_{1}^{d+2}&0&-2&(d+2)&2(d+2)&0\\ t_{1}\otimes v_{1}^{2}B_{1}^{d+1}&0&0&-4&4&0\\ t_{1}^{2}\otimes B_{2}B_{1}^{d}&-2&0&0&0&2\\ t_{1}^{2}\otimes B_{1}^{d+2}&0&-2&0&0&(d+2)\\ t_{1}^{2}\otimes v_{1}B_{1}^{d+1}&0&0&-2&0&-2\\ t_{2}\otimes B_{1}^{d+1}&0&0&-1&3&2\\ t_{1}^{3}\otimes B_{1}^{d+1}&0&0&0&3&3\end{array}$
We find that $E_{2}^{2,2d+6}$ is always zero, and the Adams chart for $d$ even is:
@@
The non-trivial $d_{3}$ differential can be observed by looking at the map of spectra $S^{\infty+2d}\rightarrow\overline{\mathbf{MTU}}(d)$ given by the generator of $\pi_{2d}(\overline{\mathbf{MTU}}(d))$. There may be an additional differential killing $E_{2}^{2,2d+8}$ but since this element is in odd degree, we will not need to know it. The Adams-Novikov spectral sequence for $d$ odd is trivial in this range. ∎
@@
Since $\pi_{2d+6}(\overline{\mathbf{MTU}}(d))$ is free abelian, there is no possible target for a torsion obstruction to $3$ complex sections. Similarly, there is no possible target for a torsion obstruction to $2$ complex sections when $d$ is odd. In the case $d$ is even, there could be an obstruction to $2$ complex sections. However, we recall from [8] that $S^{\infty+2d}$ splits off $\mathbf{MTU}(d)$ for $d$ even. The element in $E_{2}^{3,2d+3}$ comes from this split $S^{\infty+2d}$ (specifically, it can be identified with $h_{1}^{3}$), and so cannot be the target of a differential. ∎
@@
The group $\pi_{6}^{s}(BU(d-3))$ is torsion free. The torsion element in $\pi_{6}^{s}(S^{0})$ is $h_{2}\cdot h_{2}$. The map $\pi_{2d}^{s}(S^{2d-6})\rightarrow\pi_{2d}(\mathbf{MTU}(d-3))$ is a $\pi_{*}^{s}(S^{0})$ module map. In particular, it must take $h_{2}$ to 00 if $d$ is odd and we conclude that $h_{2}\cdot h_{2}\mapsto h_{2}\cdot 0=0$. Thus the torsion obstruction to $M$ having $4$ complex sections vanishes. ∎
@@
There are no non-zero differentials originating in the bottom row of the Adams-Novikov spectral sequence for $\mathbf{MTU}(1)\cong\Sigma^{\infty-2}\mathbb{C}P^{\infty}$.
@@
Then for any set of multiplicative generators $M^{2d}\in\Omega_{2d}^{U}$, we can write the cobordism class as $N^{2d}=\tilde{a}_{d}M^{2d}+\textnormal{decomposables}$